OpenStudy (anonymous):

An Irate patient complained that the cost of a doctor's visit was too high. She randomly surveyed n=20 other patients and found that the mean amount of money they spent on each doctor's visit was x=$44.80. The standard deviation of the sample was s=$3.53. Find the 95% confidence interval of the population mean. Assume the variable is normally distributed.

OpenStudy (amistre64):

zscore = (sample mean - claim)/ modified standard deviation i believe

OpenStudy (amistre64):

claim is vague but it was just cause we went over hypot testing :)

OpenStudy (amistre64):

2.575; 1.96; and 1.645 i recall; 95% has a zscore of 1.96

OpenStudy (amistre64):

E = z(mu)/(sd/sqrt(n))

OpenStudy (amistre64):

or is it mu^2?

OpenStudy (amistre64):

mu +- E is the confidence interval

OpenStudy (anonymous):

(44.8-mu)/(3.53/sqrt(20))) = z-score.. now rearranging you get mu = 44.8\[\pm\] Z-score(3.53)/sqrt(20)...i believe the z-score is 1.96. so the confidence interval is [43.25 ; 46.35]

OpenStudy (amistre64):

just to recall: n = (z o~/E)^2 sqrt(n) = z o~/E E = z o~/sqrt(n) ; yeah, that got it lol

OpenStudy (anonymous):

The confidence interval is 43.25

OpenStudy (amistre64):

1.96 ; z 3.53 ; o~ ------- 588 980 588 ------ 7.9188 hmmm, the thing is are we sure we use the zscore and not the student t?

OpenStudy (anonymous):

im pretty sure the answer i gave is right. confidence interval [43.25 ; 46.35]

OpenStudy (amistre64):

im sure it is too; just curious if it fits the z or t tho .. since n<30

OpenStudy (amistre64):

if sd known use z ; got it :)

OpenStudy (anonymous):

u cud be right,the t is used if pop SD is unknown, and we are given sample SD and not pop SD. but the fact that n>30 only comes into play if u wish to approximate one distribution using another

OpenStudy (anonymous):

t = (sample mean-mu)/(sample SD/sqrt(n))

OpenStudy (amistre64):

the t-interval gets me: (43.15, 46.45) on my ti83

OpenStudy (amistre64):

and (43.25,46.35) on the z-interval